The solution containing 4.0 g of PVC in one litre of dioxane was found to have O.P. of 6.0 × 10−4 atm. at 300 K. The mol. mass of polymer is


        (A)    6.4 × 102


        (B)    3.0 × 102


        (C)    1.6 × 105


(D)    5.6 × 104


Grade:upto college level

1 Answers

Deepak Patra
10 years ago

W2 = 4g

 

        V = 1000 mL = 11

 

         = 6.0 × 10−4 atm

 

        T = 300 K

 

        R = 0.0821 It atm K−1 mol−1

 

       

 

        M2V = W2RT

 

       

 

       

 

        = 16.42 × 104

 

        M2 = 1.6 × 105